OpenStudy (anonymous):

Subtract. Show all of your work. Be sure your answer is simplified -5/10 - 5/10

OpenStudy (anonymous):

firts change the subtrahen to addition and the sign of the minuen to negative, so that would be -5/10 + -5/10=-10/10=-1

OpenStudy (anonymous):

hope it would help friend..

OpenStudy (anonymous):

okay why add the (+)?